OpenStudy (anonymous):

A polyhedron has 5 faces and 8 edges. What is the number of vertices of the polyhedron? 5 11 8 13

OpenStudy (anonymous):

13

OpenStudy (anonymous):

can yu explain please

OpenStudy (asnaseer):

You need to use Eulers formula

OpenStudy (anonymous):

can sum one walk me through it please

OpenStudy (anonymous):

the answer is 5 thats what i got bc V+F-E=2 V+8-5=2 8-5+2=V 3+2=V 5=V

OpenStudy (asnaseer):

you have the faces/edges reversed

OpenStudy (anonymous):

oh . ok so its V+5-8=2 5-8+2=V -3+2=V -1=V

OpenStudy (anonymous):

im confused nw can i please get a explanation

OpenStudy (asnaseer):

V + 5 - 8 = 2 V - 3 = 2 <--- (5 - 8 = -3) V = 2 + 3 <--- (add 3 to both sides)

OpenStudy (anonymous):

im still a little lost

OpenStudy (asnaseer):

where abouts?

OpenStudy (asnaseer):

is it the algebra that is confusing you?

OpenStudy (anonymous):

sorta yes bc im in geometry and i dnt really remember the alegra parts anymore

OpenStudy (asnaseer):

ok, lets do this one step at a time...

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

Eulers formula tells us that for a convex polyhedron, we have: V + F - E = 2 where: V = number of vertices F = number of faces = 5 (in this case) E = number of edges = 8 (in this case)

OpenStudy (asnaseer):

ok so far?

OpenStudy (anonymous):

yess

OpenStudy (asnaseer):

ok, what we need to find is the number of vertices V, so we could first re-arrange the equation as follows: V + F - E = 2 V + F - E + E = 2 + E (adding E to both sides) V + F = 2 + E (since -E + E cancel each other out) V + F - F = 2 + E - F (subtracting F from both sides) V = 2 + E - F (since F - F cancel each other out) ok so far?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

ok, good, so we have ended up with: V = 2 + E - F now lets plug in the values E=8, F=5 to get: V = 2 + 8 - 5 = 10 - 5 = 5

OpenStudy (asnaseer):

hopefully it all makes more sense now?

OpenStudy (anonymous):

yes very much thank yuu soo much .

OpenStudy (asnaseer):

yw - I'm glad I was able to help :)